HUBBLE SPACE TELESCOPE DAILY REPORT #2192 PERIOD COVERED: 0000Z (UTC) 08/20/98 - 0000Z (UTC) 08/21/98 Daily Status Report as of 233/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.4 Completed Four Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.5 Completed Four Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.7 Completed NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.10 Completed NIC/1/2/3 7824 (Deep NIC Images of 20 High-Z Damped Ly-Alpha Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high-Z damped Lyman-Alpha (DLA) galaxy H150013. The HST imaging together with ground-based spectroscopy will unify absorption-line studies and faint galaxy imaging studies, providing a comprehensive picture of the history of assembly, star formation, gas consumption, and chemical enrichment in the DLA galaxies. The observations were completed as planned, and no problems were reported. 1.11 Completed Three Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned with no problems. 1.12 Completed Three Sets of STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.13 Completed WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet star WR69 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/1/2 7218 (Nearby Starburst Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the nearby starbursting galaxy NGC-253. This proposal encompasses a study of both the stars and interstellar material in starburst nuclei which will test ideas about the triggering of starbursts and whether starbursts propagate outwards in a nucleus. The observations were completed as planned, and no anomalies were reported. 1.15 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.16 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star clusters which are located within the Large Magellanic Cloud in order to study their formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. Near the beginning of this proposal, during a ZOE, the loss of lock described in 2.1 and HSTAR 6330 occurred, temporarily taking down the take data flag. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 Per HSTAR 6630, there was a loss of lock at 232/2244236Z that affected the proposal described in 1.16. A new #44 command was issued and lock was re-established. 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The SSR EDAC error counter was cleared twice using ROP SR-1. The NCC IP test was completed at 232/1803Z with overall success. During the test a NSSC-1 status buffer dump was performed. It completed nominally, but the first eight messages had invalid time tags. HSTAR 6629 was written. Also the PACOR-II 1.024 Mbps science dump reported errors which have since been attributed to the PACOR-II PTB board misconfiguration. A total of 2556 commands were transmitted during the test. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Later this evening, from 234/0002Z until 234/0229Z, a solar eclipse will occur. Operations personnel have ensured that Eclipse management procedures are already in place. Continuation of normal science observations and calibrations. /DMH